# On the Power of the Weisfeiler-Leman Test for Graph Motif Parameters

### v1

@@ 2 Preliminaries
where $\psi$ is a formula consisting of positive atoms, conjunction and disjunction, and $\psi^{*}$ is a conjunction of inequalities and possibly negated atoms. We call such formulas positive formulas with free constraints. We write #$\#\varphi(G)$ for the number of assignments to $\bar{x}$ for which the formula is satisfied in labeled graph $G$ (Curticapean et al., 2017).
Theorem 1 (Dell et al. (2019)).
Let $\varphi$ be a positive formula with free constraints. Then the function #$\#\varphi$ is a labeled graph motif parameter [^1] Dell et al. (2019) both additionally require $\psi$ to be in CNF/DNF. For our purposes the potential blowup incurred by transformation into CNF is of no consequence..
@@
To further illustrate the theorem, we consider a more ad-hoc parameter. Let us think we are interested in $k$-independent sets $I$ where all vertices in $I$ share a neighbor of certain type, say with vertex label either $U_{a}$ or $U_{b}$. It is difficult to intuit a priori whether this is a graph motif parameter. However, it can easily observed from the following formula that this is in fact the case:
$\phi^{\star}_{\rm IS}\ :=\ \left(\bigwedge_{1\leq i\neq j\leq k}\neg E(x_{i},x_{j})\right)\,\land\,\exists y\,\left(U_{a}(y)\lor U_{b}(y)\land\bigwedge_{i=1}^{k}E(y,x_{i})\right)\ \ \ \ \qed$
Counting graph motifs with the WL test.
@@ 3 Characterizing the WL-dimension of graph motif patterns
We now discuss our proof of Theorem 2. The full argument requires us to extend various existing results in the field to the setting of labeled graphs. We defer those proofs to the Appendix.
Let $\mathcal{F}$ be a class of labeled graphs. We write $G\equiv_{\mathcal{F}}H$ if for every $F\in\mathcal{F}$ it holds that ${\sf homs}(F,G)={\sf homs}(F,H)$. Following Roberson (2022), we define the labeled homomorphism-distinguishing closure of $\cal F$, denoted $\mathsf{homclosure}(\mathcal{F})$, as the set of all labeled graphs $L$ for which the following holds for every labeled graphs $G,H$:
@@
[…]
@@ 4 Counting occurrences of graph motif patterns
Let $\Gamma$ be a labeled graph motif parameter and suppose that the maximum treewidth of a labeled graph in ${\sf Supp}(\Gamma)$ is at most $k$. Also, let $\mathcal{C}^{k}$ denote the set of possible colors produced by the $k$WL test. Then there exists a computable function $\theta_{\Gamma}\colon\mathcal{C}^{k}\to\mathbb{Q}$ such that $\Gamma(G)=\sum_{\bar{v}\in V(G)^{k}}\theta_{\Gamma}(c^{k}(\bar{v}))$, for every labeled graph $G$.
@@
Let $F$ be a labeled graph with treewidth at most $k$ and let $\mathcal{C}^{k}$ denote the set of possible colors produced by the $k$WL test. There exists a computable function $\eta_{F}\colon\mathcal{C}^{k}\to\mathbb{N}$ such that ${\sf homs}(F,G)=\sum_{\bar{v}\in V(G)^{k}}\eta_{F}(c^{k}(\bar{v}))$, for each labeled graph $G$.
@@ 5 Determining the WL-dimension for subgraph counting
We finally move on the recognizability problem for WL-dimension. That is, for given graph motif parameter $\Gamma$, we want to know the WL-dimension of $\Gamma$. Arvind et al. (2020) previously raised the question for which labeled graph patterns $H$, the graph parameter $\mathsf{Sub}_{H}$ can be expressed by $2$WL. Here, we give a polynomial time algorithm that decides the WL-dimension for all subgraph counting problems, thus also resolving this question as a special case. For the analogous problem $\mathsf{Ind}_{H}$, we have already seen in Example 2 that the recognition problem is, in fact, trivial. In addition, we determine the WL-dimension of counting counting $k$-graphlets, i.e., all connected induced subgraphs on $k$ vertices, as illustrative examples of how further such results can be derived from the deep body of work in graph motif parameters.
We will begin with the case of subgraph counting. Recall that a minor of a labeled graph $H$ is a labeled graph $H^{\prime}$ that can be obtained from $H$ by removing nodes or edges, or contracting edges (that is, removing an edge and simultaneously merging its endpoints). By Theorem 2, and the following discussion on $\textsf{spasm}(H)$, recognising the WL-dimension of $\mathsf{Sub}_{H}$ is precisely the problem of recognising the maximum treewidth of of the homomorphic images of $H$. We can use classic results in the field of graph minors to express this as a property checkable in monadic second-order logic (MSO), for which model checking is tractable in our setting. We give a more detailed sketch of the argument below, full details are provided in the appendix.
@@
We now move on to study the WL-dimension of counting $k$-graphlets, a popular problem in graph mining and analysis of social networks Bressan et al. (2017); Jin et al. (2018). Here, no dedicated algorithm is necessary as the WL-dimension will be immediate from observations about the respective support when viewed as graph motif parameters. While the problem of counting $k$-graphlets is itself popular, its analysis in our context serves as an examples of how the recently established theory of graph motif parameters can answer these types of recognizability questions.
@@
Let $\mathsf{Ind}^{\textsc{C}}_{k}$ be graph parameter that counts the number of induced subgraphs on $k$ vertices that are connected, i.e., the number of $k$-graphlets. For graph property $\phi$, let $E^{\phi}_{k}$ be the set of all edge-subsets of the complete graph on $k$ vertices, such that the corresponding subgraph satisfies property $\phi$. Roth & Schmitt (2020) showed that the problem of counting all induced subgraphs on $k$ vertices that satisfy $\phi$ is a graph motif parameter, and that the $k$-clique is in the support of the parameter precisely when
@@
as the properties are complementary and the sums thus form an alternating sum of binomials. Since the left-hand sum is non-zero, so is the right-hand sum for property $\mathsf{C}$. We can therefore conclude that the $k$-clique is in the support of $\mathsf{Ind}^{\textsc{C}}_{k}$ [^3] This observation was already mentioned implicitly but without details in Roth (2019)., demonstrating that the maximal treewidth in the support is at least $k-1$. The argument of Roth & Schmitt (2020) also implicitly shows by construction that the support contains no graphs with more than $k$ vertices, thus also confirming that the treewidth in $\mathsf{Supp}(\mathsf{Ind}^{\textsc{C}}_{k})$ is no higher than $k-1$. Applying Theorem 2 completes the argument. ∎
Moreover, Theorem 1 is constructive and in many cases it is feasible to determine the support from logical formulations of parameters as in Example 1. For example, following the construction of Dell et al. (2019) for the formula $\phi_{IS}$ reveals that counting $k$-independent sets also has WL-dimension $k-1$. What is particularly noteworhty in this context, is that we can use the exact same methods that have been used to study the parameterized complexity of these problems. Just as the WL-dimension, the complexity of computing graph motif parameter $\Gamma$ depends precisely on the maximum treewidth in $\mathsf{Supp}(\Gamma)$ (Curticapean et al., 2017). Analysis of the complexity thus revolves around the exact same question of how high the treewidth of the basis can become.
@@ Appendix A Proof of Lemma 7
Suppose the statement were false. Then there is an edge $\{y,z\}$ with $z$ in $B(\alpha)$ but not in $B(T^{\prime})$, and $y\not\in B(\alpha)$. Let $T^{*}$ be the subtree of $T\setminus T^{\prime}$ that contains $\alpha$. By connectedness, $z$ occurs only in bags of $T^{*}$, whereas $y$ occurs in none of them. Hence the edge $\{y,z\}$ cannot be subset of any bag, contradicting that $(T,B)$ is a tree decomposition. ∎
@@
In the base case $T$ contains only $r$. Since $c^{k}(\bar{v})=c^{k}(\bar{w})$ we have that $\{v_{i}\mapsto w_{i}\}_{i\in[k]}$ is an isomorphism from $G[\bar{v}]$ to $H[\bar{w}]$. Thus $\bar{a}\mapsto\bar{v}\in\mathsf{Hom}(F[T],G)$ iff $\bar{a}\mapsto\bar{w}\in\mathsf{Hom}(F[T],H)$.
@@
If $\alpha$ is not a split node, then let $\beta$ be the leaf of $T^{\prime}$ that is adjacent (in $T^{*}$) to $\alpha$. Suppose $\alpha$ is a forget node. By assumption any $\mu\in\mathsf{Hom}(F[T^{\prime}],G)[\bar{a}\mapsto\bar{v}]$ and $\iota(\mu)$ are in colorful strict $T^{\prime}$ leaf agreement. Since $B(T)=B(T^{\prime})$, the sets of homomorphisms do not change $T^{\prime}$ and $T$. It is therefore sufficient to check that the colorful strict leaf agreement constraints are also satisfied for leaf $B(\alpha)$. In that case, $\iota$ satisfies the desired properties also for $T$. Since $B(\alpha)\subseteq B(\beta)$,any $\mu,\nu$ that are in $(\bar{d},\bar{e})$-strict $B(\beta)$-agreement, are also in $(\bar{d},\bar{e})$-strict $B(\alpha)$-agreement. It remains to verify the color constraint on $\bar{d},\bar{e}$. If $|B(\beta)|\leq k$, then we already have $c^{k}_{i}(\bar{d})=c^{k}_{i}(\bar{e})$ by assumption and are done. Otherwise we only know that $\mathsf{ct}(t,i,\bar{d}^{\ominus})=\mathsf{ct}(u,i,\bar{e}^{\ominus})$ for $\bar{d}=\bar{d}^{\ominus}t,\bar{e}=\bar{e}^{\ominus}u$. Let $z$ be the vertex from $B(\beta)$ that is forgotten by $\alpha$. Let $j$ be the index such that $\mu(z)=d_{j}$. We distinguish two cases: first, suppose that there is only one such index and there is at least one other $y\in B(\alpha)$ such that $\mu(y)=d_{j}$. Then there must be some index $j^{\prime}$ such that $d_{j^{\prime}}$ is not in $\mu(B(\alpha))$. By strict agreement the same must hold for $e_{j^{\prime}}$ and $\nu(B(\alpha))$. Hence, $\mu,\nu$ are also in $(\bar{e}[t/j^{\prime}],\bar{d}[u/j^{\prime}])$-strict $B(\alpha)$-agreement, and we have $c^{k}_{i-1}(\bar{e}[t/j^{\prime}])=c^{k}_{i-1}(\bar{d}[u/j^{\prime}])$ by equality of the color tuples. For the other case, if $z$ is the only vertex in $B(\alpha)$ that $\mu$ maps to $d_{j}$ or there is more than one index that equals $\mu(z)$, then simply take $j^{\prime}=j$ and proceed as in the other case. This concludes the case where $\alpha$ is a forget node.
@@
Let $\iota^{\prime}$ be the disjoint union of all the individual $\iota^{\prime}_{\mu}$ for all $\mu\in\mathsf{Hom}(F[T^{\prime}],G)[\bar{a}\mapsto\bar{v}]$. To see that the $\iota^{\prime}_{\mu}$ are in fact disjoint it is enough to observe that the sets $Hom(F[T],G)[\mu]$ are disjoint for different choices of $\mu$ as $dom(\mu)\subseteq V(F[T])$. Since all $\iota^{\prime}_{\mu}$ are injective, so is $\iota^{\prime}$. To see that $\iota^{\prime}$ is also surjective, suppose a $\nu\in\mathsf{Hom}(F[T],H)[\bar{a}\mapsto\bar{w}]$ not in the image of $\iota^{\prime}$. This means that $\nu$ cannot be in any set $Hom(F[T],H)[\iota(\mu)]$ for $\mu\in\mathsf{Hom}(F[T^{\prime}],G)[\bar{a}\mapsto\bar{v}]$. But $\iota$ is surjective and therefore this would mean that the projection of $\nu$ to $V(F[T^{\prime}])$ is not a homomorphism from $F[T^{\prime}$] to $H$ (extending $\bar{a}\mapsto\bar{w}$). But homomorphisms are closed under projection to induced subgraphs, so this is impossible. It follows that $\iota^{\prime}$ is also surjective and therefore is the desired bijection. ∎
As an alternative proof strategy, the main ideas from above can be used to define an algorithm that computes $Hom(F,G)[\bar{a}\mapsto\bar{v}]$ from $c^{k}(\bar{v})$. However, in practice the colors in the $k$WL test are not maintained in their explicit form as nestings of tuples and multisets but rather by compactly representing the $\mathsf{ct}$ tuples via some hash function. In that setting, the viability of such an algorithm would be limited.

### v3

@@ 2 Preliminaries
where $\psi$ is a formula consisting of positive atoms, conjunction and disjunction, and $\psi^{*}$ is a conjunction of inequalities and possibly negated atoms. We call such formulas positive formulas with free constraints. We write #$\#\varphi(G)$ for the number of assignments to $\bar{x}$ for which the formula is satisfied in labeled graph $G$. It is known that the function #$\#\varphi$ for such formulas can be expressed as a finite linear combination of homomorphism counts with projection. Göbel et al. (2024) have recently initiated the study of how homomorphism counts with projection relate to the $k$WL test. In this paper, we are content with a more restricted fragment for which no projection in the homomorphism counts is required.
Proposition 1 (Chen & Mengel (2016),Dell et al. (2019)).
Let $\varphi$ be a quantifier-free positive formula with free constraints. Then the function #$\#\varphi$ is a labeled graph motif parameter [^1] Dell et al. (2019) additionally require $\psi$ to be in CNF/DNF. For our purposes the potential blowup incurred by transformation into CNF is of no consequence..
@@
Counting graph motifs parameters and the WL test.
@@ 3 Characterizing the WL-dimension of graph motif patterns
We now discuss our proof of Theorem 2. The full argument requires us to extend various existing results in the field to the setting of labeled graphs. We defer those proofs to the appendix. Let $\mathcal{F}$ be a class of labeled graphs. We write $G\equiv_{\mathcal{F}}H$ if for every $F\in\mathcal{F}$ it holds that ${\sf homs}(F,G)={\sf homs}(F,H)$. Following Roberson (2022), we define the labeled homomorphism-distinguishing closure of $\cal F$, denoted $\mathsf{homclosure}(\mathcal{F})$, as the set of all labeled graphs $L$ for which the following holds for every labeled graphs $G,H$:
@@
For all labeled graphs $G,H$ we have $G\equiv_{k\text{WL}}H$ if and only if $G\equiv_{\mathcal{LT}_{k}}H$.
@@ 4 Counting occurrences of graph motif patterns
Let $\Gamma$ be a labeled graph motif parameter and suppose that the maximum treewidth of a labeled graph in ${\sf Supp}(\Gamma)$ is at most $k$. Also, let $\mathcal{C}^{k}$ denote the set of possible colors produced by the $k$WL test. Then there exists a function $\theta_{\Gamma}\colon\mathcal{C}^{k}\to\mathbb{Q}$ such that $\Gamma(G)=\sum_{\bar{v}\in V(G)^{k}}\theta_{\Gamma}(c^{k}(\bar{v}))$, for every labeled graph $G$.
@@
Let $F$ be a labeled graph with treewidth at most $k$ and let $\mathcal{C}^{k}$ denote the set of possible colors produced by the $k$WL test. There exists a function $\eta_{F}\colon\mathcal{C}^{k}\to\mathbb{N}$ such that ${\sf homs}(F,G)=\sum_{\bar{v}\in V(G)^{k}}\eta_{F}(c^{k}(\bar{v}))$, for each labeled graph $G$.
@@ 5 Determining the WL-dimension for subgraph counting
We finally move on the recognizability problem for WL-dimension. That is, for given graph motif parameter $\Gamma$, we want to know the WL-dimension of $\Gamma$. Arvind et al. (2020) previously raised the question for which labeled graph patterns $H$, the graph parameter $\mathsf{Sub}_{H}$ can be expressed by $2$WL. Here, we give a polynomial time algorithm that decides the WL-dimension for all subgraph counting problems, thus also resolving this question as a special case. For the analogous problem $\mathsf{Ind}_{H}$, we have already seen in Example 2 that the recognition problem is, in fact, trivial. In addition, we determine the WL-dimension of counting $k$-graphlets, i.e., all connected induced subgraphs on $k$ vertices, as illustrative examples of how further such results can be derived from the deep body of work in graph motif parameters.
We will begin with the case of subgraph counting. Recall that a minor of a labeled graph $H$ is a labeled graph $H^{\prime}$ that can be obtained from $H$ by removing nodes or edges, or contracting edges (that is, removing an edge and simultaneously merging its endpoints). By Theorem 2, and the following discussion on $\textsf{spasm}(H)$, recognising the WL-dimension of $\mathsf{Sub}_{H}$ is precisely the problem of recognising the maximum treewidth of the homomorphic images of $H$. We can use classic results in the field of graph minors to express this as a property checkable in monadic second-order logic (MSO), for which model checking is tractable in our setting. We give a more detailed sketch of the argument below, full details are provided in the appendix.
@@
We now move on to study the WL-dimension of counting $k$-graphlets, a popular problem in graph mining and analysis of social networks Bressan et al. (2017); Jin et al. (2018). Here, no dedicated algorithm is necessary as the WL-dimension will be immediate from observations about the respective support when viewed as graph motif parameters. While the problem of counting $k$-graphlets is itself popular, its analysis in our context serves as a example of how the recently established theory of graph motif parameters can answer these types of recognizability questions.
@@
Let $\mathsf{Ind}^{\textsf{C}}_{k}$ be graph parameter that counts the number of induced subgraphs on $k$ vertices that are connected, i.e., the number of $k$-graphlets. For graph property $\phi$, let $E^{\phi}_{k}$ be the set of all edge-subsets of the complete graph on $k$ vertices, such that the corresponding subgraph satisfies property $\phi$. Roth & Schmitt (2020) showed that the problem of counting all induced subgraphs on $k$ vertices that satisfy $\phi$ is a graph motif parameter, and that the $k$-clique is in the support of the parameter precisely when
@@
as the properties are complementary and the sums thus form an alternating sum of binomials. Since the left-hand sum is non-zero, so is the right-hand sum for property $\mathsf{C}$. We can therefore conclude that the $k$-clique is in the support of $\mathsf{Ind}^{\textsf{C}}_{k}$ [^3] This observation was already made implicitly but without details in Roth (2019)., demonstrating that the maximal treewidth in the support is at least $k-1$. The argument of Roth & Schmitt (2020) also implicitly shows by construction that the support contains no graphs with more than $k$ vertices, thus also confirming that the treewidth in $\mathsf{Supp}(\mathsf{Ind}^{\textsf{C}}_{k})$ is no higher than $k-1$. Applying Theorem 2 completes the argument. ∎
Moreover, Proposition 1 is constructive and in many cases it is feasible to determine the support from logical formulations of parameters as in Example 1. For example, following the construction of Dell et al. (2019) for the formula $\phi_{IS}$ reveals that counting $k$-independent sets also has WL-dimension $k-1$. What is particularly noteworthy in this context, is that we can use the exact same methods that have been used to study the parameterized complexity of these problems. Just as the WL-dimension, the complexity of computing graph motif parameter $\Gamma$ depends precisely on the maximum treewidth in $\mathsf{Supp}(\Gamma)$ (Curticapean et al., 2017). That is, analysis of the complexity of computing a graph motif parameter and analysis of the WL-dimension of the parameter revolve around the same question of how high the treewidth of the basis can become.
@@ Appendix A Proof of Lemma 7
Suppose the statement were false. Then there is an edge $\{y,z\}$ with $z$ in $B(\alpha)$ but not in $B(T^{\prime})$, and $y\not\in B(\alpha)$. Let $T^{*}$ be the subtree of $T\setminus T^{\prime}$ that contains $\alpha$. By connectedness, $z$ occurs only in bags of $T^{*}$, whereas $y$ occurs in none of them. Hence, the edge $\{y,z\}$ cannot be subset of any bag, contradicting that $(T,B)$ is a tree decomposition. ∎
@@
In the base case $T$ contains only $r$. Since $c^{k}(\bar{v})=c^{k}(\bar{w})$ we have that $\{v_{i}\mapsto w_{i}\}_{i\in[k]}$ is an isomorphism from $G[\bar{v}]$ to $H[\bar{w}]$. Thus, $\bar{a}\mapsto\bar{v}\in\mathsf{Hom}(F[T],G)$ iff $\bar{a}\mapsto\bar{w}\in\mathsf{Hom}(F[T],H)$.
@@
If $\alpha$ is not a split node, then let $\beta$ be the leaf of $T^{\prime}$ that is adjacent (in $T^{*}$) to $\alpha$. Suppose $\alpha$ is a forget node. By assumption any $\mu\in\mathsf{Hom}(F[T^{\prime}],G)[\bar{a}\mapsto\bar{v}]$ and $\iota(\mu)$ are in colorful strict $T^{\prime}$ leaf agreement. Since $B(T)=B(T^{\prime})$, the sets of homomorphisms do not change $T^{\prime}$ and $T$. It is therefore sufficient to check that the colorful strict leaf agreement constraints are also satisfied for leaf $B(\alpha)$. In that case, $\iota$ satisfies the desired properties also for $T$. Since $B(\alpha)\subseteq B(\beta)$, any $\mu,\nu$ that are in $(\bar{d},\bar{e})$-strict $B(\beta)$-agreement, are also in $(\bar{d},\bar{e})$-strict $B(\alpha)$-agreement. It remains to verify the color constraint on $\bar{d},\bar{e}$. If $|B(\beta)|\leq k$, then we already have $c^{k}_{i}(\bar{d})=c^{k}_{i}(\bar{e})$ by assumption and are done. Otherwise, we only know that $\mathsf{ct}(t,i,\bar{d}^{\ominus})=\mathsf{ct}(u,i,\bar{e}^{\ominus})$ for $\bar{d}=\bar{d}^{\ominus}t,\bar{e}=\bar{e}^{\ominus}u$. Let $z$ be the vertex from $B(\beta)$ that is forgotten by $\alpha$. Let $j$ be the index such that $\mu(z)=d_{j}$. We distinguish two cases: first, suppose that there is only one such index and there is at least one other $y\in B(\alpha)$ such that $\mu(y)=d_{j}$. Then there must be some index $j^{\prime}$ such that $d_{j^{\prime}}$ is not in $\mu(B(\alpha))$. By strict agreement the same must hold for $e_{j^{\prime}}$ and $\nu(B(\alpha))$. Hence, $\mu,\nu$ are also in $(\bar{e}[t/j^{\prime}],\bar{d}[u/j^{\prime}])$-strict $B(\alpha)$-agreement, and we have $c^{k}_{i-1}(\bar{e}[t/j^{\prime}])=c^{k}_{i-1}(\bar{d}[u/j^{\prime}])$ by equality of the color tuples. For the other case, if $z$ is the only vertex in $B(\alpha)$ that $\mu$ maps to $d_{j}$ or there is more than one index that equals $\mu(z)$, then simply take $j^{\prime}=j$ and proceed as in the other case. This concludes the case where $\alpha$ is a forget node.
@@
Let $\iota^{\prime}$ be the disjoint union of all the individual $\iota^{\prime}_{\mu}$ for all $\mu\in\mathsf{Hom}(F[T^{\prime}],G)[\bar{a}\mapsto\bar{v}]$. To see that the $\iota^{\prime}_{\mu}$ are in fact disjoint it is enough to observe that the sets $\mathsf{Hom}(F[T],G)[\mu]$ are disjoint for different choices of $\mu$ as $dom(\mu)\subseteq V(F[T])$. Since all $\iota^{\prime}_{\mu}$ are injective, so is $\iota^{\prime}$. To see that $\iota^{\prime}$ is also surjective, suppose a $\nu\in\mathsf{Hom}(F[T],H)[\bar{a}\mapsto\bar{w}]$ not in the image of $\iota^{\prime}$. This means that $\nu$ cannot be in any set $\mathsf{Hom}(F[T],H)[\iota(\mu)]$ for $\mu\in\mathsf{Hom}(F[T^{\prime}],G)[\bar{a}\mapsto\bar{v}]$. But $\iota$ is surjective and therefore this would mean that the projection of $\nu$ to $V(F[T^{\prime}])$ is not a homomorphism from $F[T^{\prime}$] to $H$ (extending $\bar{a}\mapsto\bar{w}$). But homomorphisms are closed under projection to induced subgraphs, so this is impossible. It follows that $\iota^{\prime}$ is also surjective and therefore is the desired bijection. ∎
As an alternative proof strategy, the main ideas from above can be used to define an algorithm that computes $\mathsf{Hom}(F,G)[\bar{a}\mapsto\bar{v}]$ from $c^{k}(\bar{v})$. An argument along this direction also demonstrate computability of the functions in Theorem 6 and Lemma 7 in the general case (in the uniform case where $\mathcal{C}^{k}$ is finite, it suffices to enumerate graphs until all colors in $\mathcal{C}^{k}$ are observed). However, in practice the colors in the $k$WL test are not maintained in their explicit form as nestings of tuples and multisets but rather by compactly representing the $\mathsf{ct}$ tuples via some hash function. In that setting, the viability of such an algorithm would be limited.